Find the equation of a circle when the 2 end points of a diameter is given by A(-2, 1) , B(-1, 2)
1
Expert's answer
2021-02-02T04:30:01-0500

The center of the circle has the coordinates: "O(-\frac32,\frac32)". It is the midpoint of the segment A(-2,1), B(-1,2). The radius of the circle is: "r=\sqrt{(-\frac32+2)^2+(\frac32-1)^2}=\sqrt{\frac12}=\frac{\sqrt{2}}{2}" . The equation is "(x+\frac32)^2+(y-\frac32)^2=\frac12" .

We can also consider the problem with complex numbers. I.e., we may present points in the following form : A:-2+i; B:-1+2i and O: "-\frac32+\frac32i". The radius is "r=\frac{\sqrt{2}}{2}" (the distance between O and A). The equation of the circle is: "|z-z_0|^2=\frac12" , where "z" presents a complex variable. I.e., "z=x+iy" and "z_0=-\frac32+\frac32i"